youdointlikeme:

Which of the following fractions is equivalent to 10/12 ?

Imagine:

What are the following fractions?

youdointlikeme:

40/80 20/30 4/5 12/14

lowkey:

Let's simplify the 10/12

lowkey:

Do you know how to?

youdointlikeme:

no

lowkey:

Wait hold on

youdointlikeme:

ok

lowkey:

None of these fractions is equivalent...

umm:

You can find fractions equivalent to 10/12 by multiplying the numerator and denominator by the same whole number.

Imagine:

12/14

Imagine:

Because 10+2=12/12+2=14

youdointlikeme:

thx you get a meadial

Imagine:

In that case 12/14 is correct

RodeWolf:

Yes I agree

Imagine:

@youdointlikeme Thank you!

umm:

\[\frac{ 10 }{ 12 }=\frac{ 10 x2}{ 12x2 }=\frac{ 10x3 }{12x3 }=\frac{ 10x4 }{ 12x4 }\] \[\frac{ 10 }{ 12 }=\frac{ 20 }{ 24 }=\frac{ 30 }{ 36 }=\frac{ 40 }{ 48 }\]

umm:

Supposed to be 4 groups.